a truck rolling down a hill 500ft with coefficient of friction 0.02 and gradient 2.5, what is the speed of the truck
1
Expert's answer
2020-10-05T15:31:42-0400

As per the given question,

Height of the inclined plane "(h)=500ft"

Coefficient of friction "(\mu)=0.02"

Gradient "(m)=\tan\theta=2.5"

"\Rightarrow \frac{height}{base}=2.5"


"\Rightarrow \frac{500}{base}=2.5"

"\Rightarrow base=\dfrac{500}{2.5}ft"


"\Rightarrow base=\frac{152.4}{2.5}m"

"\Rightarrow base=60.96m"

"\sin \theta=\frac{152.4}{164.13}=0.928"


"\cos\theta =\frac{60.96}{164.13}=0.37"

Let the resultant acceleration of the truck is "=a"

So, "a=g\sin\theta-\mu g\cos\theta"

"\Rightarrow a=9.8\times 0.92-0.02\times 9.8\times 0.37 m/sec^2"

"\Rightarrow a=8.943 m/sec^2"

"\Rightarrow v^2=u^2+2as"

"\Rightarrow v=\sqrt{2as}"

"\Rightarrow v=\sqrt{2\times 8.94\times 164.13}"

"\Rightarrow v=54.17m/sec"